OpenStudy (anonymous):

medal will be rewarded for whoever ha correct Algebra 1 answer

OpenStudy (mendicant_bias):

Is it clear enough that you should assume n = 0, or n = 1 for the start of 3?

OpenStudy (mendicant_bias):

No, because 3^0 = 1. So 3^(0)*5 = 5. Your initial digit would definitely be five.

OpenStudy (mendicant_bias):

Yes. (Sorry, didn't intend to confuse you.) As for number four: |-1|^2 is still just one. So f(2) = 2*|-1|^2 = 2*(1)^2 = 2*1 = 2.

OpenStudy (mendicant_bias):

And as a general rule, one put to any power will always be one. Your first answer is correct; it's a geometric sequence. In your second one, even if n = 1, none of the answers properly align to that; This would be the set resulting from either n = 0 or n = 1:\[n _{0} = 5*3^{0} = 5*1 = 5\]\[n _{1} = 5*3^{1} = 5*3 = 15\]\[n _{2} = 5*3^{2} = 5*9 = 45\]\[n _{3} = 5*3^{3} = 5*27 = 135\]|||-|||-|||-|||

OpenStudy (mendicant_bias):

Yes, two and three are correct, and four should be \[(2)*|-1|^{2} = 2*(1)^{2} = 2*1 = 2\]

OpenStudy (anonymous):

thank you again for explaining! Can you help me one one more please?

OpenStudy (mendicant_bias):

No problem, and sure.

OpenStudy (anonymous):

*one more

OpenStudy (mendicant_bias):

Okay, so since you have a set of points, you can do a few things to test this: You ca find the slope of the line that may or may not be produced, and test whether the slope remains the same. I'm sure you know the slope formula. \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]From here, just plug in the coordinates \[(-2, -7)|||(-1, 1)|||(0, 8)|||(1, 17)|||(2, 15)\]...\[\frac{ (-7)-(1) }{ (-2) - (-1) } = \frac{ -8 }{ -1 } = 8\] Now test if this is the same as slopes between other points: ...\[\frac{ (1)-(8) }{ (-1)-(0) }= \frac{ -7 }{ -1 } = 7\] A linear function should have a slope that is consistent regardless of where you sample it. Does this function have the same apparent slope?

OpenStudy (anonymous):

So far, yes

OpenStudy (mendicant_bias):

Are you sure? Look at the results I got. Are those the same?

OpenStudy (mendicant_bias):

Exactly!

OpenStudy (anonymous):

thank you!!!!

OpenStudy (mendicant_bias):

Glad I could help. If you have any other questions, just ask. I'm sure that plenty of people will be able to help.